Susan is 10 1/3 yrs. old while her father is 25 3/4 yrs. older than Susan. how age is Susan's father?
Answer by josmiceli(19441)

10 1/3 + 25 3/4 = 10 4/12 + 25 9/12
------------------------------------
10 4/12 + 25 9/12 = 35 13/12
----------------------------------
35 13/12 = 36 1/12
-------------------
Note that 1/12 of a year is 1 month, so
Susan's father is 36 yrs 1 month old